The four choices for the last event of the semester (our "final"). There are great NOVA videos out there and we will choose one for our final grade. The titles/topics of the episodes are:

Hunting the Hidden Dimension (Fractals) F

The Great Math Mystery (the usefulness of math in the real world) M

Inside Einstein's Mind (his thought experiments) E

Origami REvolution (the uses of origami in science) O

1. Please rank your choice of topic from your 1st choice to 4th choice using the letters above.

For example, if your first choice is Math Mystery, then Fractals, Einstein, then Origami, your would vote

1st M

2nd F

3rd E

4th O

As you know from the reading, if you don't rank order all 4, your vote will not be used.

2. Using what we learned in the unit on probability and counting, how many different ballots can there be for these 4 choices (how many ways can the be ordered?) Show your work.

Type your ballot (preferences) and the answer to part 2 in the text box on Canvas.
Answer #1

1) 1st M

2nd E

3rd O

4th F

2) The total number of different ballots for these choices is 24

[ for the 1st choice we can choose any 1 topic from 4 topics in 4C1(=4) ways, for 2nd choice we choose 1 topic from the remaining 3 topics in 3C1 (=3) ways, for the 3rd choice we choose 1 topic from the remaining 2 topics in 2C1 (=2) ways and we selaect the last remaining topic as pur 4th choice. Thus total number of different choices= 4C1*3C1*2C1=4*3*2=24 ]